Suppose in a hypothetical economy that velocity is 5, the money supply is $5,000, Real GDP is 2500 units of output, and the price level is $10. If the money supply doubled over a short time period to 10,000, the sample quantity theory would predict that..
A) real GDP would double to 5,000 units
B) Velocity will decline dramatically such that there will be little change in either out put or price level
C) the price level will fall to $5
D) The price level will double
1
Expert's answer
2014-12-03T11:25:49-0500
V = 5, M = $5,000, Real GDP Y = 2500 units of output, P = $10. If the money supply doubled over a short time period to 10,000, the sample quantity theory would predict that M*V = P*Y, so if M1 = 2M, then P*Y will double too. But as output can be changed only in the long run, the price level will double. So, the right answer is D) The price level will double.